OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

when two line are parallel. their slopes are equal, now you have y = -3x + 12 here the slope is -3 now the slope of the required line is also -3 now you have slope = m = -3 and point = (-1,6) use the following formula to find the equation y-y1 = m(x-x1) here m = -3 and x1 = -1 and y1 = 6 do it now